# RAGSys: Item-Cold-Start Recommender as RAG System

### v1

@@ Abstract.
Information Retrieval, Search, Recommender Systems, Large Language Models, Few-Shot Learning, In-Context-Learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: IR-RAG @ SIGIR24; July 28, 2024; Washington, DCisbn: 978-1-XXXX-XXXX-X/XX/XXccs: Information systems Information extractionccs: Computing methodologies Natural language generationccs: Information systems Recommender systems
1. Introduction
@@
Large Language Models (LLMs) have emerged as a powerful tool for natural language processing, demonstrating remarkable abilities in areas like text completion, summarization, and question answering. One of their most intriguing capabilities is their potential to learn ”common sense” – general knowledge about the world that allows them to reason and make inferences beyond the literal meaning of text. This has fueled excitement about the possibility of achieving zero-shot learning, where LLMs can solve unseen problems without any prior training on specific tasks.
@@
However, a crucial distinction exists between generic public knowledge and the specific private knowledge required for most real-world use cases. While LLMs excel at generic text completion or chat-like interactions, practical applications often demand solving specific and repeatable problems within a particular domain. This typically necessitates knowledge specific to a business or organization, such as understanding internal processes, up-to-date product details, or customer behavior.
@@
Fine-tuning, a technique where LLMs are trained on large datasets tailored to the target task, offers a path towards adapting LLMs to these domain-specific needs. Yet, fine-tuning presents significant challenges. When trained on tasks-specific data, LLMs tend to forget knowledge and skills gained in the initial training, a phenomenon referred to as Catastrophic Forgetting (Luo et al., 2023). Consequently, a fine-tuned LLM loses some of its ability to generalize to novel examples that aren’t well represented in its fine-tuning training data. Moreover, while fine-tuning allows an LLM to memorize task-specific information, it doesn’t necessarily allow the LLM to reason about that information (Berglund et al., 2023). As a final consideration, keeping LLMs constantly up-to-date using fine-tuning can be infeasible, especially for domains with frequently changing information like e-commerce product inventory, whereas it is easy to update a database in real-time from which information is retrieved.
@@
As an alternative to fine-tuning, In-Context Learning (ICL) offers a promising approach for leveraging LLMs in scenarios with limited data. This approach exploits the demonstrated ability of LLMs for ”meta-learning” – essentially, learning how to learn. In (Bai et al., 2023), the authors prove the capacity of LLMs to effectively ingest in-context training data points and solve statistical optimization problems such as gradient descent. ICL enables practitioners to leverage Retrieval Augmented Generation (RAG), that is enriching the input prompt by information that is retrieved in real-time (Ram et al., 2023). We refer to (Dong et al., 2023) for a recent survey on ICL.
@@
This paper focuses on few-shot learning and the retrieval of relevant demonstrations for this process, where a demonstration is some text which is included in the LLM’s context to demonstrate how the LLM should formulate correct answers. Few-shot learning presents a well-structured problem, allowing us to evaluate the quality of the retrieval algorithm using established classification metrics. Crucially, it offers a powerful method to achieve fine-tuning-like behavior, steering the output of the LLM towards the desired outcome even with limited data. Interestingly, increasing the context size in prompts beyond a certain point yields diminishing returns. The most impactful information resides within a relatively small set of well-chosen demonstrations, rather than overloading the prompt with vast amounts of data (Machlab and Battle, 2024). This highlights the importance of effective retrieval strategies, transforming $k$-shot learning into a top-$k$ information retrieval problem at its core.
@@
Building upon this concept, this paper identifies desirable properties for a RAG system under the framework of few-shot learning. We demonstrate that state-of-the-art retrieval systems in this context resemble item-cold-start recommender systems. Unlike exact search algorithms that prioritize precision and recall, our focus is on discovery, by maximizing the set of collective information gain from the retrieved demonstrations. This necessitates solving various trade-offs between query relevance, quality scoring, as well as diversity algorithms to ensure a variety of informative examples are surfaced. Furthermore, we propose a method for evaluating RAG system performance through the subsequent performance of the enriched LLM on established NLP tasks like question answering or text generation. This methodology offers a valuable approach to directly assessing diversity and quality-based retrieval systems, which removes the need to define a subjective diversity score, a historically challenging aspect of evaluating such systems in academic settings (Clarke et al., 2008).
@@
To summarize, in this paper we study the impact of diversity and quality bias in retrieving demonstrations for ICL. We start by reviewing the use of diversity and other biases in both ICL and Information Retrieval works. We then propose a method for evaluating the performance of different retrieval algorithms. The method measures the impact of providing a context to an LLM by calculating the probability than an LLM generates a correct answer given the provided context. Then we present experiments and results demonstrating the impact of diversity on an LLM’s ability to generate correct answers. Finally we discuss the applicability of state-of-the-art ICL retrieval algorithms in real-world setting, and show that recommendation engines offer a better solution than semantic search engines.
2. Related Work
@@
2.1. Few-Shot In-Context Learning with Retrieval Augmented Generation
@@
Within the context of RAG, few-shot learning can be defined as a specific scenario where the ”documents” retrieved are actually ”examples” used to guide the LLM. These examples can also be referred to interchangeably as ”In-Context Examples” (ICE) or ”demonstrations”. The importance of ICL in achieving state-of-the-art LLM performance is undeniable, with its ubiquitous presence in top benchmarks across various domains. Consequently, ICL research is a rapidly evolving field with numerous proposed algorithms.
2.1.1. Pure Diversity
@@
Several noteworthy ICL approaches have emerged that address the challenge of retrieving informative examples for few-shot learning. Some methods only promote the diversity of the demonstrations, like in (Yu et al., 2023) where the authors utilize $k$-means clustering in a dense embeddings space to achieve diversity. By applying $k$-means to the sentence embeddings of the demonstrations, this approach ensures that the retrieved examples cover a variety of semantic spaces, inherently increasing the mutual information of the retrieved set, but without taking into account the relevancy to the query.
2.1.2. Pure Quality
@@
Other approaches focuses on identifying examples where the LLM exhibits low token-level uncertainty. In (Gonen et al., 2023) the authors analyze token probabilities within candidate 0-shot prompts. By prioritizing prompts where the LLM has the highest generation likelihood (low perplexity), this approach aims to select examples that hold the potential for significant learning gains for the LLM. The intuition that the authors give is that a prompt that is more expected by the LLM is more likely to help it extracting the relevant information. Accessing the per-token probabilities for all examples incurs a significant compute cost, but can be pre-computed as they do not depend on the query.
2.1.3. Pure Relevance
@@
Notably, a connection can be drawn between traditional full-text search algorithms and pure relevance approaches. In (Agrawal et al., 2023) the authors use BM25 (Robertson et al., 1995), a well-established retrieval ranking function commonly used in information retrieval tasks. This approach essentially leverages the strengths of BM25 in identifying examples with terms highly similar to the query, to select the most relevant examples for the specific task at hand. This strategy ensures the retrieved examples are topically relevant to the task while potentially introducing some variation in the specific phrasing or wording used.
@@
Finally, neural ranking, one of the most commonly used ICL approach, typically yielding superior results (Rubin et al., 2022; Wang et al., 2022; Li et al., 2023; Ye et al., 2023b; Wang et al., 2024), is maximizing similarity in a dense embeddings space. These methods, like KATE (Liu et al., 2022a), utilize $k$-Nearest Neighbors (kNN) search using the cosine distance of sentence embeddings to retrieve the examples most semantically similar to the prompt. Scaling this method leverages vector search algorithms, commonly used in large-scale information retrieval tasks, where efficient retrieval of semantically similar documents is crucial.
@@
While general purpose pre-trained embeddings like BERT (Devlin et al., 2019) form a strong baseline, learning specific embeddings for retrieval, and in particular ICL, is a very active area of research. In (Khattab and Zaharia, 2020) the authors build upon BERT and introduce ColBERT that improves the retrieval performances by an order of magnitude. Other embeddings model have been proposed for ICL retrieval in (Wang et al., 2022) and (Li et al., 2023). Some authors also explored training full language models (Rubin et al., 2022), as well as LLM (Wang et al., 2024), showing further improvements compared to traditional embedding-based approaches. While leading to superior results, these supervised neural ranking models for learning-to-rank necessitate orders of magnitude more training examples data, that is typically not available to practitioners. In addition, without any explicit metric space such as dense embeddings, efficient retrieval indexing such as (Beygelzimer et al., 2006) cannot be used.
2.1.4. Diversity/Relevance Trade-off
@@
While both relevance and diversity are crucial for effective few-shot learning, methods yielding the best ICL results combine these two paradigms rather than prioritizing one over the other. This is achieved by maximizing a carefully balanced trade-off between semantic similarity to the prompt and diversity of the retrieved examples. Unsupervised techniques can be adapted to prioritize the selection of examples that are both relevant to the prompt and dissimilar to each other. In (Levy et al., 2023) the authors introduce a greedy method to select relevant demonstrations while ensuring enough coverage. They define specific coverage strategies adapted to the problem of program generation. In (Su et al., 2023) the authors employs an active learning setting, where a voting algorithm selects a set of examples penalizing the top-$k$ closest from already selected ones, using cosine distance in an embedding space.
@@
The most popular unsupervised approach for achieving this balance between relevance and diversity is Maximal Marginal Relevance (MMR). MMR retrieves a set of examples by iteratively selecting the example that maximizes a linear combination of the scores of relevance to the prompt and dissimilarity to the previously retrieved examples. It was analyzed in (Ye et al., 2023a) for ICL and was shown to outperform simpler methods. Alternatively to MMR, Determinantal Point Processes (DPP) has been used in (Ye et al., 2023b) to optimize the joint information of the selected $k$ examples. However, exactly solving the DPP optimization being NP-hard, hence the authors also employs greedy maximization.
2.2. Diversity in Information Retrieval and Online Learning
@@ 2.2.1. In Information Retrieval
The MMR algorithm was analyzed in (Carbonell and Goldstein, 1998), and compared against other approached like KL-divergence optimization in (Zhai et al., 2003). Pure retrieval algorithms typically optimize for recall, not information maximization. Agreeing on a diversity objective function remains a challenge. Diversity is sometimes introduced as a heuristic to cover possible interpretations of the same query, instead of minimizing information overlap from near-duplicated results. In (Clarke et al., 2008) the authors leverage a concept of information nuggets with documents to estimate the redundancy of the set of retrieved documents. Topic modeling is also employed, such as (Agrawal et al., 2009) that uses a taxonomy of categories labelling the documents and the queries. The desired properties of diverse retrieval are furthermore characterized in (Gollapudi and Sharma, 2009). A various set of similarity methods and diversification algorithms are analyzed in (Vieira et al., 2011) on sparse features vectors. Among diversity evaluation methods based on topic modelling, three notable criteria used in the TREC Diversity track (Clarke et al., 2009a), ERR-IA (Chapelle et al., 2009), $\alpha$-nDCG@$k$ (Clarke et al., 2008), and NRBP (Clarke et al., 2009b), are compared in (Clarke et al., 2011).
2.2.2. In Recommender Systems
@@
Within IR, the recommender system literature brings an additional point-of-view on studying diversity in retrieval, by focusing on the benefit of diverse results for a user, instead of evaluating methods against a potentially arbitrary relevance/diversity trade-off. The difficulty of evaluating the impact of diversity, and the necessity for large scale real-world recommendation studies has been explored in (Ziegler et al., 2005). In (Vargas and Castells, 2011) and (Vargas, 2014) the authors model the user behavior conditioned on the set of retrieved items. In (Zheng et al., 2021) the authors improve the diversity versus relevance trade-off in recommender systems by directly learning a ranking model that favor diversity, instead of only applying diversity re-ranking methods.
2.2.3. In Online Learning
@@
Learning a trade-off between relevancy and diversity also naturally occurs in related online frameworks such as active learning, multi-armed bandits and Bayesian optimization. In (Radlinski et al., 2008) the authors modify a learning-to-rank algorithm from users feedback, to inherently learn diverse rankings and demonstrate a positive impact on the original relevance metric. Other approaches such as (Zhu et al., 2014) also introduce diversity in learning-to-rank algorithms while preserving the offline settings, but then are limited to evaluate using direct diversity measures.
@@
Within batch-mode Bayesian optimization, in (Desautels et al., 2014) and (Contal et al., 2013) the authors analyze two greedy exploration/exploitation algorithms to select the next batch of items maximizing the cumulative reward. Like with recommender systems, these online settings exemplify the theoretical and empirical importance of diversifying the selected set of items despite the true objective function only including pure relevance, the cumulative reward.
2.3. Quality Biases in Information Retrieval
@@
2.3.1. Content-Based Quality Biases
Content-based methods leverage existing signals inside the documents themselves to identify potentially lower-quality content. Examples include spam detection scores developed in works like (Ntoulas et al., 2006) and (Cormack et al., 2011). By incorporating such scores during retrieval, the system can prioritize higher quality documents. More sophisticated content-based approaches don’t limit at spam classification, but extract more generic quality features the content of documents. Works like (Bendersky et al., 2011) explore features such as stop-word statistics or entropy of the documents to generate quality scores. The authors demonstrate that biasing standard retrieval using these features leads to improved retrieval effectiveness even using unbiased IR metrics like nDCG.
2.3.2. Graph-Based Quality Biases
Instead of relying on the content itself, graph-based algorithms inherently capture implicit quality signals within their ranking model. PageRank, a seminal algorithm for web search ranking introduced in (Brin and Page, 1998), exemplifies this approach. PageRank leverages the links structure between web articles to assign higher importance to web pages that are linked to by other high-quality pages. This process implicitly prioritizes documents with a higher perceived quality based on the quality of their in-links.
2.3.3. Connections to Recommender Systems

### v2

# Information Retrieval’s Role in RAG Systems (IR-RAG) - 2024

@@ Abstract
keywords:
[…]
1 Introduction
@@
Large Language Models (LLMs) have emerged as a powerful tool for natural language processing, demonstrating remarkable abilities in areas like text completion, summarization, and question answering [1]. One of their most intriguing capabilities is their potential to learn "common sense" – general knowledge about the world that allows them to reason and make inferences beyond the literal meaning of text. This has fueled excitement about the possibility of achieving zero-shot learning, where LLMs can solve unseen problems without any prior training on specific tasks [2].
@@
However, a crucial distinction exists between generic public knowledge and the specific private knowledge required for most real-world use cases. While LLMs excel at generic text completion or chat-like interactions, practical applications often demand solving specific and repeatable downstream tasks within a particular domain [3]. This typically necessitates knowledge specific to a business or organization, such as understanding internal processes, up-to-date product details, or customer behavior.
@@
Fine-tuning, a technique where LLMs are trained on large datasets tailored to the target task, offers a path towards adapting LLMs to these domain-specific needs. Yet, fine-tuning presents significant challenges. When trained on tasks-specific data, LLMs tend to forget knowledge and skills gained in the initial training, a phenomenon referred to as Catastrophic Forgetting [4]. Consequently, a fine-tuned LLM loses some of its ability to generalize to novel examples that aren’t well represented in its fine-tuning training data. Moreover, while fine-tuning allows an LLM to memorize task-specific information, it doesn’t necessarily allow the LLM to reason about that information [5]. As a final consideration, keeping LLMs constantly up-to-date using fine-tuning can be infeasible, especially for domains with frequently changing information like e-commerce product inventory, whereas it is easy to update a database in real-time from which information is retrieved.
@@
As an alternative to fine-tuning, In-Context Learning (ICL) offers a promising approach for leveraging LLMs in scenarios with limited data. This approach exploits the demonstrated ability of LLMs for "meta-learning" – essentially, learning how to learn. In [6], the authors prove the capacity of LLMs to effectively ingest in-context training data points and solve statistical optimization problems such as gradient descent. ICL enables practitioners to leverage Retrieval Augmented Generation (RAG), that is enriching the input prompt by information that is retrieved in real-time [7]. We refer to [8] for a recent survey on ICL.
@@
This paper focuses on few-shot learning and the retrieval of relevant demonstrations for this process, where a demonstration is some text which is included in the LLM’s context to demonstrate how the LLM should formulate correct answers. Few-shot learning presents a well-structured problem, allowing us to evaluate the quality of the retrieval algorithm using established classification metrics. Crucially, we show that enriching a language model with a few-shot example retriever offers a powerful method to achieve fine-tuning-like behavior, steering the output of the LLM towards the desired outcome even with limited data. Interestingly, increasing the context size in prompts beyond a certain point yields diminishing returns. The most impactful information resides within a relatively small set of well-chosen demonstrations, rather than overloading the prompt with vast amounts of data [9]. This highlights the importance of effective retrieval strategies, transforming $k$-shot learning into a top-$k$ information retrieval problem at its core.
@@
Building upon this concept, this paper identifies desirable properties for a RAG system under the framework of few-shot learning. We demonstrate that state-of-the-art retrieval systems in this context resemble item-cold-start recommender systems. Unlike exact search algorithms that prioritize precision and recall, our focus is on discovery, by maximizing the set of collective information gain from the retrieved demonstrations. This necessitates solving various trade-offs between query relevance, quality scoring, as well as diversity algorithms to ensure a variety of informative examples are surfaced. Furthermore, we propose a method for evaluating RAG system performance through the subsequent performance of the enriched LLM on established NLP tasks like question answering or text generation. This methodology offers a valuable approach to directly assessing diversity and quality-based retrieval systems, which removes the need to define a subjective diversity score, a historically challenging aspect of evaluating such systems in academic settings [10].
@@
To summarize, in this paper we study the impact of diversity and quality bias in retrieving demonstrations for ICL. We start by reviewing the use of diversity and other biases in both ICL and Information Retrieval works. We then propose a method for evaluating the performance of different retrieval algorithms. Then we present experiments and results demonstrating the impact of diversity and quality bias on an LLM’s ability to generate correct answers. Finally we discuss the applicability of state-of-the-art ICL retrieval algorithms in real-world setting, and show that recommendation engines offer a better solution than semantic search engines.
2 Related Work
@@
2.1 Few-Shot In-Context Learning with Retrieval Augmented Generation
@@
Within the context of RAG, few-shot learning can be defined as a specific scenario where the "documents" retrieved are actually "examples" used to guide the LLM. These examples can also be referred to interchangeably as "In-Context Examples" (ICE) or "demonstrations". The importance of ICL in achieving state-of-the-art LLM performance is undeniable, with its ubiquitous presence in top benchmarks across various domains. Consequently, ICL research is a rapidly evolving field with numerous proposed algorithms.
2.1.1 Pure Diversity
@@
Several noteworthy ICL approaches have emerged that address the challenge of retrieving informative examples for few-shot learning. Some methods only promote the diversity of the demonstrations, like in [11] where the authors utilize $k$-means clustering in a dense embeddings space to achieve diversity. By applying $k$-means to the sentence embeddings of the demonstrations, this approach ensures that the retrieved examples cover a variety of semantic spaces, inherently increasing the mutual information of the retrieved set, but without taking into account the relevancy to the query.
2.1.2 Pure Quality
@@
Other approaches focuses on identifying examples where the LLM exhibits low token-level uncertainty. In [12] the authors analyze token probabilities within candidate 0-shot prompts. By prioritizing prompts where the LLM has the highest generation likelihood (low perplexity), this approach aims to select examples that hold the potential for significant learning gains for the LLM. The intuition that the authors give is that a prompt that is more expected by the LLM is more likely to help it extracting the relevant information. Accessing the per-token probabilities for all examples incurs a significant compute cost, but can be pre-computed as they do not depend on the query.
2.1.3 Pure Relevance
@@
Notably, a connection can be drawn between traditional full-text search algorithms and pure relevance approaches. In [13] the authors use BM25 [14], a well-established retrieval ranking function commonly used in information retrieval tasks. This approach essentially leverages the strengths of BM25 in identifying examples with terms highly similar to the query, to select the most relevant examples for the specific task at hand. This strategy ensures the retrieved examples are topically relevant to the task while potentially introducing some variation in the specific phrasing or wording used.
@@
Finally, neural ranking, one of the most commonly used ICL approach, typically yielding superior results [15, 16, 17, 18, 19], is maximizing similarity in a dense embeddings space. These methods, like KATE [20], utilize $k$-Nearest Neighbors (kNN) search using the cosine distance of sentence embeddings to retrieve the examples most semantically similar to the prompt. Scaling this method leverages vector search algorithms, commonly used in large-scale information retrieval tasks, where efficient retrieval of semantically similar documents is crucial.
@@
While general purpose pre-trained embeddings like BERT [21] form a strong baseline, learning specific embeddings for retrieval, and in particular ICL, is a very active area of research. In [22] the authors build upon BERT and introduce ColBERT that improves the retrieval performances by an order of magnitude. Other embeddings model have been proposed for ICL retrieval in [16] and [17]. Some authors also explored training full language models [15], as well as LLM [19], showing further improvements compared to traditional embedding-based approaches. While leading to superior results, these supervised neural ranking models for learning-to-rank necessitate orders of magnitude more training examples data, that is typically not available to practitioners. In addition, without any explicit metric space such as dense embeddings, efficient retrieval indexing such as [23] cannot be used.
2.1.4 Diversity/Relevance Trade-off
@@
While both relevance and diversity are crucial for effective few-shot learning, methods yielding the best ICL results combine these two paradigms rather than prioritizing one over the other. This is achieved by maximizing a carefully balanced trade-off between semantic similarity to the prompt and diversity of the retrieved examples. Unsupervised techniques can be adapted to prioritize the selection of examples that are both relevant to the prompt and dissimilar to each other. In [24] the authors introduce a greedy method to select relevant demonstrations while ensuring enough coverage. They define specific coverage strategies adapted to the problem of program generation. In [25] the authors employs an active learning setting, where a voting algorithm selects a set of examples penalizing the top-$k$ closest from already selected ones, using cosine distance in an embedding space.
@@
The most popular unsupervised approach for achieving this balance between relevance and diversity is Maximal Marginal Relevance (MMR). MMR retrieves a set of examples by iteratively selecting the example that maximizes a linear combination of the scores of relevance to the prompt and dissimilarity to the previously retrieved examples. It was analyzed in [26] for ICL and was shown to outperform simpler methods. Alternatively to MMR, Determinantal Point Processes (DPP) has been used in [18] to optimize the joint information of the selected $k$ examples. However, exactly solving the DPP optimization being NP-hard, hence the authors also employs greedy maximization.
2.2 Diversity in Information Retrieval and Online Learning
@@ 2.2.1 In Information Retrieval
The MMR algorithm was analyzed in [27], and compared against other approached like KL-divergence optimization in [28]. Pure retrieval algorithms typically optimize for recall, not information maximization. Agreeing on a diversity objective function remains a challenge. Diversity is sometimes introduced as a heuristic to cover possible interpretations of the same query, instead of minimizing information overlap from near-duplicated results. In [10] the authors leverage a concept of information nuggets with documents to estimate the redundancy of the set of retrieved documents. Topic modeling is also employed, such as [29] that uses a taxonomy of categories labelling the documents and the queries. The desired properties of diverse retrieval are furthermore characterized in [30]. A various set of similarity methods and diversification algorithms are analyzed in [31] on sparse features vectors. Among diversity evaluation methods based on topic modelling, three notable criteria used in the TREC Diversity track [32], ERR-IA [33], $\alpha$-nDCG@$k$ [10], and NRBP [34], are compared in [35].
2.2.2 In Recommender Systems
@@
Within IR, the recommender system literature brings an additional point-of-view on studying diversity in retrieval, by focusing on the benefit of diverse results for a user, instead of evaluating methods against a potentially arbitrary relevance/diversity trade-off. The difficulty of evaluating the impact of diversity, and the necessity for large scale real-world recommendation studies has been explored in [36]. In [37] and [38] the authors model the user behavior conditioned on the set of retrieved items. In [39] the authors improve the diversity versus relevance trade-off in recommender systems by directly learning a ranking model that favor diversity, instead of only applying diversity re-ranking methods.
2.2.3 In Online Learning
@@
Learning a trade-off between relevancy and diversity also naturally occurs in related online frameworks such as active learning, multi-armed bandits and Bayesian optimization. In [40] the authors modify a learning-to-rank algorithm from users feedback, to inherently learn diverse rankings and demonstrate a positive impact on the original relevance metric. Other approaches such as [41] also introduce diversity in learning-to-rank algorithms while preserving the offline settings, but then are limited to evaluate using direct diversity measures.
@@
Within batch-mode Bayesian optimization, in [42] and [43] the authors analyze two greedy exploration/exploitation algorithms to select the next batch of items maximizing the cumulative reward. Like with recommender systems, these online settings exemplify the theoretical and empirical importance of diversifying the selected set of items despite the true objective function only including pure relevance, the cumulative reward.
2.3 Quality Biases in Information Retrieval
@@
2.3.1 Content-Based Quality Biases
Content-based methods leverage existing signals inside the documents themselves to identify potentially lower-quality content. Examples include spam detection scores developed in works like [44] and [45]. By incorporating such scores during retrieval, the system can prioritize higher quality documents. More sophisticated content-based approaches don’t limit at spam classification, but extract more generic quality features the content of documents. Works like [46] explore features such as stop-word statistics or entropy of the documents to generate quality scores. The authors demonstrate that biasing standard retrieval using these features leads to improved retrieval effectiveness even using unbiased IR metrics like nDCG.
2.3.2 Graph-Based Quality Biases
Instead of relying on the content itself, graph-based algorithms inherently capture implicit quality signals within their ranking model. PageRank, a seminal algorithm for web search ranking introduced in [47], exemplifies this approach. PageRank leverages the links structure between web articles to assign higher importance to web pages that are linked to by other high-quality pages. This process implicitly prioritizes documents with a higher perceived quality based on the quality of their in-links.
2.3.3 Connections to Recommender Systems